The weights for newborn babies is approximately normally distributed with a mean of 6.4 pounds and a standard deviation of 1.2 pounds.

Consider a group of 1600 newborn babies:

1. How many would you expect to weigh between 3 and 9 pounds?

2. How many would you expect to weigh less than 8 pounds?

3. How many would you expect to weigh more than 6 pounds?

4. How many would you expect to weigh between 6.4 and 10 pounds?
Answer #1

solution/ given values : 1.2 p(3-6.4 <xc9-6.4) = 6.4, standard deviation Men soup of 1600 new born babies =P(-2.83€ 242.17) =PloL243) 10-6.4) 4-6.4 LXL 1-2) 4 P (64<x< 10) : Pikach <<< =P[2<3) - Pezzo) = 0.9987 -0.5 0.4987 -0.4987x1600 797.92